OpenStudy (anonymous):

Check my work please? Q: What is the mass of 1.8 × 1021 molecules of H2SO4? A:(1.8x10^21 Molecules)(1 Mole of H2SO4 / 6.02x10^23) (130 G of H2SO4/ 1 Mole of H2SO4) = .39 Grams of H2SO4 Check my work please? Q: What is the mass of 1.8 × 1021 molecules of H2SO4? A:(1.8x10^21 Molecules)(1 Mole of H2SO4 / 6.02x10^23) (130 G of H2SO4/ 1 Mole of H2SO4) = .39 Grams of H2SO4 @Chemistry

OpenStudy (anonymous):

\[1.8x10^21 \left(\begin{matrix} 1 Mole of H2SO4\\ 6.02x10^23 Molecules\end{matrix}\right)\left(\begin{matrix}130.14 Grams of H2SO4 \\ 1 Mole of H2SO4\end{matrix}\right) = .39 Grams of H2SO4\] That should be a bit clearer...

OpenStudy (anonymous):

BINGO